# Cleaning Up the Streets: Understanding Motivations, Mental Models, and Concerns of Users Flagging Social Media Posts

### v2

@@ 3.1. Participant Recruitment
Our university’s IRB [^1] We will specify the University name after the anonymous peer review is completed. approved this study on Jan 31, 2023. We recruited participants by posting on Twitter, NextDoor, and Craigslist. In selecting these platforms to place our recruitment calls on, we aimed to reach out to a diverse set of potential participants. We asked candidates to submit an online form that included questions about whether they use social media daily, which social media sites they use, whether they have encountered toxic content on social media, whether they flagged a post in the past week and their demographic information. We included an option to attach a screenshot for a post the users recently flagged. We also included an open-ended question: “What is your perspective on how social media platforms can improve flagging mechanisms?”
We reviewed submissions and screened for responses with legitimate and thoughtful answers. In selecting interview candidates, we drew inspiration from prior literature that advocates for researchers and practitioners to give voice to the perspectives of users who are more likely to be vulnerable to online harm  (Duggan, 2017; Musgrave et al., 2022; Blackwell et al., 2017) as well as literature showing such users are more likely to be exposed to content that is harmful (Duggan, 2017; Jhaver et al., 2018b). Therefore, we prioritized offering interviews to candidates who self-identified as belonging to marginalized groups, such as Black and LGBTQ+ users. Table 2 shows the demographic details of our participants.

### v3

@@ 3.1. Participant Recruitment
Our university’s IRB [^1] We will specify the University name after the anonymous peer review is completed. approved this study on Jan 31, 2023. We recruited participants by posting on Twitter, NextDoor, and Craigslist. In selecting these platforms to place our recruitment calls on, we aimed to reach out to a diverse set of potential participants. We asked candidates to submit an online form that included questions about whether they use social media daily, which social media sites they use, whether they have encountered toxic content on social media, whether they flagged a post in the past week, and their demographic information. We included an option to attach a screenshot for a post the users recently flagged. We also included an open-ended question: “What is your perspective on how social media platforms can improve flagging mechanisms?”
We reviewed submissions and screened for responses with legitimate and thoughtful answers. In selecting interview candidates, we drew inspiration from prior literature that advocates for researchers and practitioners to give voice to the perspectives of users who are more likely to be vulnerable to online harm  (Duggan, 2017; Musgrave et al., 2022; Blackwell et al., 2017) as well as literature showing such users are more likely to be exposed to content that is harmful (Duggan, 2017; Jhaver et al., 2018b). Additionally, we note prior examples of research on user perspectives on content moderation that also oversampled users with marginalized identities  (Jhaver et al., 2023). Therefore, we prioritized offering interviews to candidates who self-identified as belonging to marginalized groups, such as Black and LGBTQ+ users. Table 1 shows the demographic details of our participants.